# Deep Learning for Melt Pool Depth Contour Prediction From Surface Thermal Images via Vision Transformers

### v2

@@ Abstract
Insufficient overlap between the melt pools produced during Laser Powder Bed Fusion (L-PBF) can lead to lack-of-fusion defects and deteriorated mechanical and fatigue performance. In-situ monitoring of the melt pool subsurface morphology requires specialized equipment that may not be readily accessible or scalable. Therefore, we introduce a machine learning framework to correlate in-situ two-color thermal images observed via high-speed color imaging to the two-dimensional profile of the melt pool cross-section. Specifically, we employ a hybrid CNN-Transformer architecture to establish a correlation between single bead off-axis thermal image sequences and melt pool cross-section contours measured via optical microscopy. In this architecture, a ResNet model embeds the spatial information contained within the thermal images to a latent vector, while a Transformer model correlates the sequence of embedded vectors to extract temporal information. Our framework is able to model the curvature of the subsurface melt pool structure, with improved performance in high energy density regimes compared to analytical melt pool models. The performance of this model is evaluated through dimensional and geometric comparisons to the corresponding experimental melt pool observations.
@@ Introduction
Laser Powder Bed Fusion (L-PBF) is a metal additive manufacturing process, which decreases material waste and increases local control when compared to conventional manufacturing processes [1, 2]. Due to these advantages, L-PBF has become a useful production technique for industrial applications requiring parts with complex internal structure [3, 4, 5, 6, 7]. During L-PBF, a product is iteratively built by fusing successive layers of metal alloy powder together, using a laser heat source to induce melting at a specific cross-section [8]. However, due to the process complexity and short length-time scales involved during L-PBF, the stochastic formation of defects can increase performance variability, limiting adoption for use cases with strict part specifications [9, 10]. For instance, the transient behavior of the instantaneous molten metal zone near laser impact with the substrate directly changes the residual stress, microstructure, and porosity within the part [11, 12, 13]. Unstable vapor cavities within the part can collapse and cause spherical gas voids, and insufficient melting between layers results in lack-of-fusion porosity [14, 15]. These defects have a significant effect on the mechanical and fatigue performance of the finalized parts, localizing stress and reducing the life cycle until fatigue induced failure [16, 17, 18].
@@
In response, researchers have developed small-scale thermal imaging tools for monitoring the behavior of the melt pool, enabling more direct linkage to the melt pool dimensions and defect regimes. For instance, Gaikwad et al. monitored the melt pool with coaxial high-speed video cameras at two different wavelengths [25]. From this data, they designed features to link the sensor information to the type of porosity present. In an alternate work, Larsen et al. present FlawNet, a method to detect anomalous melt pool instances potentially indicative of porosity through autoregressive training on the temporal dynamics of the melting process [39]. While these works show promise in detecting porosity within the finalized part, these models are built upon binary classification algorithms, and do not directly characterize the degree of porosity present in the material, or account for potential pore formation mechanisms.
Therefore, we propose an approach that links experimental, high-speed in-situ videos of the melt pool to the ex-situ melt pool morphology beneath the surface. Specifically, high-speed off-axis image sequences of the melt pool surface temperature observed during printing are collected and mapped to the corresponding solidified melt pool cross-section. The ability to predict the melt pool contour from surface images provides a pathway towards in-situ control for defect mitigation. Our model employs a transformer-based architecture, originally developed for sequence modeling in natural language processing tasks [40], to analyze these image sequences. The transformer architecture uses attention mechanisms to learn how elements in a sequence affect each other and inform downstream prediction tasks. To reduce the complexity of our model, we decompose the architecture into spatial and temporal components. Specifically, we leverage a convolutional encoder to model the spatial information, and the transformer model to extract temporal information. With this framework, we also examine the influence of transfer learning from multiphysics simulation data to the experimental domain to reduce the requirements for manual data collection. The performance of the model is evaluated by comparing the dimensions of the predicted cross-sectional areas the shape of the predicted contours, and the Intersection-over-Union (IoU) score of the contour reconstructions. Finally, we validate our model through comparisons to ex-situ characterization of multi-track prints.
@@
Figure 2: a) The melt pool surface temperature images are sampled at intervals of 44 µs, corresponding to a frame rate of 22500 frames per second. The temperature distribution changes from frame to frame. At low exposure times, more of the melt pool structure is visible, but the temperature core of the melt pool near the laser is saturated. At lower exposure times, less of the melt pool is directly observable, but the center of the melt pool is no longer saturated. b) Optical micrographs of the cross-sectioned tracks at varying energy densities. c)The melt pool depth contour profiles for varying energy densities. There are slight variations due to the uncertainty present between successive melt tracks.
@@
After the single bead scans were run, the 316L SS plates were cross-sectioned, mounted, sanded, and polished. To identify the melt pool boundary, the polished samples were electro-etched following procedure 13b in the ASTM Standard Practice for Microetching Metals and Alloys (E-407). Then, the melt pool boundary profile was extracted from microscope images of the cross-sections by drawing around the edge of the melt pool in the ImageJ software. Four single beads were processed at each of the sixty-four combinations of power and velocity with eight powers ranging from 50 W to 400 W and eight velocities ranging from 0.3 m/s to 2.4 m/s. However, not all of the melt pool cross-sections were measurable due to either small melt pool size, over-etching, or under-etching.
@@ Results
We separate the dataset into training and testing partitions based on a 75% training, 25% testing split. Specifically, 25% of the power-velocity combinations present within the dataset are not considered during training in order to evaluate the performance of the model on unseen data. The train-test split is applied in the space of power-velocity combinations to avoid data leakage in the temporal sequences.
@@
The performance of the transformer model on each configuration is listed in Table 2. Notably, we observe that the most accurate predictions are obtained with the temperature estimates derived from the ratiometric temperature images. The IoU of the predicted cross-sections appears is almost identical between an exposure time of $E=4\mu s$, and an exposure time of $E=20\mu s$. With this configuration, the trained model achieves the lowest error between the predicted and experimentally observed melt pool dimensions, and the largest IoU score.
@@
Temporal Transformer 0.764 $\pm$ 0.038 20.086 $\pm$ 3.956 9.917 $\pm$ 3.001 0.767 $\pm$ 0.012 19.234 $\pm$ 1.301 8.138 $\pm$ 1.804
ViT 0.767 $\pm$ 0.017 25.546 $\pm$ 0.347 11.230 $\pm$ 0.323 0.790 $\pm$ 0.013 24.559 $\pm$ 0.016 9.166 $\pm$ 0.735
UNet 0.763 $\pm$ 0.017 26.538 $\pm$ 3.006 11.759 $\pm$ 0.724 0.763 $\pm$ 0.055 30.471 $\pm$ 2.124 13.372 $\pm$ 1.794
@@
Figure 5: A comparison of the extracted melt pool dimensions for contour predictions on the test partition of the experimental dataset, for three different model architectures. a) The area correlation between the ground truth and predicted melt pool contours, for the U-Net, Vision Transformer, and Temporal Transformer models. The dotted lines indicate a deviation of ±$\pm 6500\;\mu m^{2}$ from the ideal prediction. b) The area correlation between the ground truth and predicted melt pool contours, for the U-Net, Vision Transformer, and Temporal Transformer models. The dotted lines indicate a deviation of ±$\pm 45\;\mu m$ from the ideal prediction.
Each model is trained for 50 epochs on the same train-test dataset split, and the pre-processing steps are held constant. More details regarding the specific configuration of each model are provided in the Appendix. We evaluate the performance of each of the three architectures for two different exposure times on the surface temperature estimates (Table 1).
We first examine the comparison between a U-Net architecture for this task, and the implemented transformer architecture. We find that our model out-performs the U-Net on the contour shape metrics, and achieves comparable accuracy on the IoU metric. Specifically, the $\pm$$1\sigma$ confidence interval of the temporal transformer network falls directly within the confidence interval of the U-Net architecture’s performance on the IoU metric. We graphically compare the performance of the three models by comparing the experimental melt pool dimensions with the dimensions of the corresponding predicted melt pool (Figure 5. When examining the model predictions on a process parameter basis, we observe that at higher energy densities, the variation within the predicted melt pool dimensions becomes more significant (Figure 5). However, the implemented transformer architecture has much less variation in the predictions produced over time by a single model. Notably, our model is able to achieve this performance with significantly smaller model with 7.5 ×$\times 10^{6}$ parameters, compared to the 8 ×$\times 10^{7}$ parameters within the implemented standard U-Net model.
@@
The approach implemented for integrating simulation data with experimental data is based upon transfer learning, where models are first trained via simulation results and subsequently fine-tuned on experimental observations. In this framework, a model will first be trained to predict the melt pool cross-sectional morphology using simulation-based surface images and cross-sections. To do so, we create a dataset of 300 SS316L single-track bare plate runs using the CFD software FLOW-3D® [54]. FLOW-3D is a multiphysics package which solves the coupled partial differential equations governing mass transfer, heat transfer, and momentum transfer within the melt pool. FLOW-3D incorporates considerations for the physical phenomena taking place during the melting process, including phase change, laser reflection, and evaporation. A more detailed description of these considerations is provided within [49, 30].
These simulations are constructed with FLOW-3D (Version 12.0, Release 7) to cover the power-velocity space with a laser power range of 130 W to a laser power of 500 W at an increment of 30 W, and a velocity range of 0.3 m/s to 1.5 m/s at an increment of 0.05 m/s.
@@
To ensure consistency with the experimental monitoring data, the input thermal images and target depth contours are rescaled to match the experimental monitoring resolution. Specifically, a $320\mu m$ by 320 $\mu m$ area around the laser is extracted and upscaled by a factor of two to obtain a 64 $\times$ 64 pixel image resolution. Further interpolation and grid rescaling is applied to ensure the spatial resolution of the simulated surface thermal images matches the 5.6 $\mu m$ pixel resolution of the experimental thermal images.
@@
To examine the learned embedding space of the model trained on simulated surface temperatures, we perform dimensionality reduction on the 128-dimensional readout token optimized during the training process. To examine how this token changes throughout the dataset in a visually intuitive manner, we apply the dimensionality reduction algorithm t-distributed Stochastic Neighbor Embedding (t-SNE), to produce a reduced order 2-dimensional vector representation for each sample in the dataset [55]. The correlation of these low-dimensional representations with the properties of the melt pool are shown in Figure 7, where each point represents the compressed read-out token associated with a specific data sample. In Figure 7a), intuitive trends are observed within the 2-dimensional latent space. Specifically, samples with similar depths in the unseen test dataset are observed to be clustered together in the latent space, and a smooth transition between the shallow melt pools and deeper melt pools is also apparent. In Figure 7b), these trends are also shown to hold for the energy density of each melt pool, also demonstrating that the model is able to learn generalizable, intuitive representations of the underlying physical mechanisms of melt pool formation.
Figure 7: Accuracy metrics after fine-tuning a model trained on FLOW-3D with experimentally collected data. The model is able to recover performance comparable to the full experimental dataset more quickly than a model trained from a random initialization. Each model instance was trained for 100 epochs on the FLOW-3D dataset, and finetuned for 10 epochs on a subset of the experimental data. a) The $R^{2}$ of the area correlation compared between the pretrained model and a randomly initialized model. b) The Intersection-over-Union score of the area correlation compared between a pretrained model and the randomly initialized model.
@@
Figure 8: a) A comparison of the Intersection-over-Union (IoU) metric as a function of the amount of data used for model training. These comparisons are made for both a model trained with randomly initialized weights, and a model that has been transfer learned from FLOW-3D simulation data. b) A comparison of the correlation between the area extracted from the predicted cross-section to the area extracted from the ground truth cross-section, a function of the amount of data used for model training. These comparisons are made for both a model trained with randomly initialized weights, and a model that has been transfer learned from FLOW-3D simulation data.
In order to apply the information learned by the model in simulation to the experimentally collected dataset, we first train the model for 100 epochs on the simulation dataset as described earlier. Next, we fine-tune the trained model on a subset of the training partition of the single-track melt pool cross-sections. To do so, we initialize the model with the weights learned from the FLOW-3D training process, and continue the training process with a randomly selected sample of the training partition of the experimental dataset. We evaluate the performance of the fine-tuned model on the original test partition of the dataset as the fraction of training data available as input is varied. In Figure 4a), the correlation between the ground truth and predicted melt pool cross-sectional areas is shown to increase dramatically when a model pre-trained on simulation data is used for prediction, compared to a network with randomly initialized weights. A similar effect is observed in Figure 4b) for the IoU score of the predicted cross-sections. Therefore, pre-training the model on the melt pool contours generated via simulation enables predictions with an area correlation of $R^{2}$ of 0.65, using only 25% of the power-velocity combinations cross-sectioned for this work.

### v3

@@ Abstract
Insufficient melt depths during Laser Powder Bed Fusion (L-PBF) can lead to lack-of-fusion defects and deteriorated mechanical and fatigue performance. In-situ monitoring of the melt pool subsurface morphology requires specialized equipment that may not be readily accessible or scalable. Therefore, we introduce a machine learning framework to correlate in-situ two-color thermal images observed via high-speed color imaging to the two-dimensional profile of the melt pool cross-section. Specifically, we employ a hybrid CNN-Transformer architecture to establish a correlation between single bead off-axis thermal image sequences and melt pool cross-section contours measured via optical microscopy. In this architecture, a ResNet model embeds the spatial information contained within the thermal images to a latent vector, while a Transformer model correlates the sequence of embedded vectors to extract temporal information. The performance of this model is evaluated through dimensional and geometric comparisons to the corresponding experimental melt pool observations. Our framework is able to model the curvature of the subsurface melt pool structure, with improved performance in high energy density regimes compared to analytical melt pool models. Additionally, we note that the use of ratiometric temperature estimates improves the accuracy of the model predictions compared to monochromatic imaging.
@@ Introduction
Laser Powder Bed Fusion (L-PBF) is a metal additive manufacturing process that decreases material waste and increases local control when compared to conventional manufacturing processes [1, 2]. Due to these advantages, L-PBF has become a useful production technique for industrial applications requiring parts with complex internal structure [3, 4, 5, 6, 7]. During L-PBF, a product is iteratively built by fusing successive layers of metal alloy powder together, using a laser heat source to induce melting at a specific cross-section [8]. However, due to the process complexity and short length and time scales involved, the stochastic formation of defects can increase performance variability and limit adoption for use cases with strict part specifications [9, 10]. For instance, the transient behavior of the instantaneous melt pool directly changes the residual stress, microstructure, and porosity within the part [11, 12, 13]. At high energy density, unstable vapor cavities within the part can collapse and cause spherical gas voids, while at low energy density, insufficient melting between layers results in lack-of-fusion porosity [14, 15]. These defects have a significant effect on the mechanical and fatigue performance of the finalized parts, localizing stress and reducing the life cycle until fatigue induced failure [16, 17, 18].
@@
In response, researchers have developed small-scale melt pool imaging tools, enabling more direct linkage to the melt pool dimensions and defect regimes. For instance, Gaikwad et al. monitored the melt pool with coaxial high-speed video cameras at two different wavelengths [25]. From this data, they designed features to link the sensor information to the type of porosity present. In an alternate work, Larsen et al. present FlawNet, a method to detect anomalous melt pool instances potentially indicative of porosity through autoregressive training on the temporal dynamics of the melting process [39]. While these works show promise in detecting porosity within the finalized part, these models are built upon binary classification algorithms, and do not directly characterize the degree of porosity present in the material, or account for potential pore formation mechanisms.
Therefore, we propose an approach that links experimental, high-speed in-situ videos of the melt pool to the ex-situ melt pool morphology beneath the surface. Specifically, high-speed off-axis image sequences of the melt pool surface temperature observed during printing are collected and mapped to the corresponding solidified melt pool cross-section. The ability to predict the melt pool contour from surface images provides a pathway towards in-situ control for defect mitigation. Our model employs a transformer-based architecture, originally developed for sequence modeling in natural language processing tasks [40], to analyze these image sequences. The transformer architecture uses attention mechanisms to learn how elements in a sequence affect each other and inform downstream prediction tasks. To reduce the complexity of our model, we decompose the architecture into spatial and temporal components. Specifically, we leverage a convolutional encoder to model the spatial information, and the transformer model to extract temporal information. With this framework, we also examine the influence of transfer learning from simulation data to the experimental domain to reduce the amount of manual data collection required. The performance of the model is evaluated by comparing the dimensions of the predicted cross-sectional areas the shape of the predicted contours, and the Intersection-over-Union (IoU) score of the contour reconstructions. Finally, we validate our model through comparisons to ex-situ characterization of multi-track prints.
@@
Figure 2: a) The melt pool surface temperature images are sampled at intervals of 44 µs, corresponding to a frame rate of 22, 500 frames per second. The temperature distribution changes from frame to frame. At high exposure times, more of the melt pool structure is visible, but the temperature core of the melt pool near the laser is saturated. At lower exposure times, less of the melt pool is directly observable, but the center of the melt pool is no longer saturated. b) Optical micrographs of the cross-sectioned tracks at varying energy densities. c)The melt pool depth contour profiles for varying energy densities. There are slight variations due to the variability present between successive melt tracks.
@@
After the single bead scans were run, the 316L SS plates were cross-sectioned, mounted, sanded, and polished. To identify the melt pool boundary, the polished samples were electro-etched following procedure 13b in the ASTM Standard Practice for Microetching Metals and Alloys (E-407). Then, the melt pool boundary profile was extracted from microscope images of the cross-sections by drawing around the edge of the melt pool in the ImageJ software. Four single beads were processed at each of the sixty-four combinations of power and velocity with eight powers ranging from 50 W to 400 W and eight velocities ranging from 0.3 m/s to 2.4 m/s. However, twenty-four melt pool cross-sections were not measurable due to either small melt pool size, over-etching, or under-etching.
@@ Results
We separate the dataset into training and testing partitions based on a 75% training, 25% testing split. Specifically, 25% of the power-velocity combinations present within the dataset are not considered during training to enable the evaluation of the performance of the model on unseen data. The train-test split is applied in the space of power-velocity combinations to avoid data leakage in the temporal sequences.
@@
The performance of the transformer model on each configuration is listed in Table 2, where the best-performing configuration is indicated in bold. Notably, we observe that the most accurate predictions are obtained with the temperature estimates derived from the ratiometric temperature images. The IoU of the predicted cross-sections appears is almost identical between an exposure time of $E=4\mu s$, and an exposure time of $E=20\mu s$. With this configuration, the trained model achieves the lowest error between the predicted and experimentally observed melt pool dimensions, and the largest IoU score.
@@
Temporal Transformer 0.76 $\pm$ 0.04 20.09 $\pm$ 3.96 9.92 $\pm$ 3.00 0.77 $\pm$ 0.01 19.23 $\pm$ 1.30 8.14 $\pm$ 1.80
ViT 0.77 $\pm$ 0.02 25.55 $\pm$ 0.35 11.23 $\pm$ 0.32 0.79 $\pm$ 0.01 24.56 $\pm$ 0.02 9.17 $\pm$ 0.74
UNet 0.76 $\pm$ 0.02 26.54 $\pm$ 3.01 11.76 $\pm$ 0.72 0.76 $\pm$ 0.06 30.47 $\pm$ 2.12 13.37 $\pm$ 1.79
@@
Figure 5: A comparison of the extracted melt pool dimensions for contour predictions on the test partition of the experimental dataset, for three different model architectures. a) The area correlation between the ground truth and predicted melt pool contours, for the U-Net, Vision Transformer, and Temporal Transformer models. The dotted lines indicate a deviation of ±$\pm 6500\;\mu m^{2}$ from the ideal prediction. b) The depth correlation between the ground truth and predicted melt pool contours, for the U-Net, Vision Transformer, and Temporal Transformer models. The dotted lines indicate a deviation of ±$\pm 45\;\mu m$ from the ideal prediction.
Each model is trained for 50 epochs on the same train-test dataset split, and the pre-processing steps are held constant. More details regarding the specific configuration of each model are provided in the Appendix. We evaluate the performance of each of the three architectures for two different exposure times on the surface temperature estimates (Table 1), where the best-performing configuration is also indicated in bold.
We first examine the comparison between a U-Net architecture for this task, and the implemented transformer architecture. We find that our model out-performs the U-Net on the contour shape metrics, and achieves comparable accuracy on the IoU metric. Specifically, the $\pm$$1\sigma$ confidence interval of the temporal transformer network falls directly within the confidence interval of the U-Net architecture’s performance on the IoU metric. We graphically compare the performance of the three models by comparing the experimental melt pool dimensions with the dimensions of the corresponding predicted melt pool (Figure 5). When examining the model predictions on a process parameter basis, we observe that at higher energy densities, the variation within the predicted melt pool dimensions becomes more significant (Figure 5). However, the implemented transformer architecture has much less variation in the predictions produced over time by a single model. Notably, our model is also significantly smaller with 7.5 ×$\times 10^{6}$ parameters compared to the 8 ×$\times 10^{7}$ parameters within the implemented standard U-Net model.
@@
The approach implemented for integrating simulation data with experimental data is based upon transfer learning, where models are first trained via simulation results and subsequently fine-tuned on experimental observations. In this framework, a model is first trained to predict the melt pool cross-sectional morphology using simulation-based surface images and cross-sections. This study is performed for both data produced from high-fidelity, multiphysics simulations of the melt pool behavior with the CFD software FLOW-3D® [54], and data produced from simplified analytical models of the heat conduction process [22]. This enables us to evaluate the need for performing more computationally expensive simulations to create data, as opposed to using less accurate low-fidelity models. To evaluate the performance of the high-fidelity simulation data, we first create a dataset of 300 SS316L single-track bare plate runs using FLOW-3D. FLOW-3D is a multiphysics package which solves the coupled partial differential equations governing mass transfer, heat transfer, and momentum transfer within the melt pool. FLOW-3D incorporates considerations for the physical phenomena taking place during the melting process, including phase change, laser reflection, and evaporation. A more detailed description of these considerations is provided within [49, 30].
These simulations are constructed with FLOW-3D (Version 12.0, Release 7) to cover the power-velocity space with a laser power range of 130 W to a laser power of 500 W at an increment of 30 W, and a velocity range of 0.3 m/s to 1.5 m/s at an increment of 0.05 m/s. The material parameters of the simulation were validated following the process described in [49].
@@
To ensure consistency with the experimental monitoring data, the input thermal images and target depth contours are rescaled to match the experimental monitoring resolution. Specifically, a $320\mu m\times 320\mu m$ area around the laser is extracted and upscaled by a factor of two to obtain a 64$\times$64 pixel image resolution. Further interpolation and grid rescaling is applied to ensure the spatial resolution of the simulated surface thermal images matches the 5.6 $\mu m$ pixel resolution of the experimental thermal images.
@@
To examine the learned embedding space of the model trained on simulated surface temperatures, we perform dimensionality reduction on the 128-dimensional readout token optimized during the training process. To examine how this token changes throughout the dataset in a visually intuitive manner, we apply the dimensionality reduction algorithm t-distributed Stochastic Neighbor Embedding (t-SNE), to produce a reduced order 2-dimensional vector representation for each sample in the dataset [55]. The correlation of these low-dimensional representations with the properties of the melt pool are shown in Figure 7, where each point represents the compressed read-out token associated with a specific data sample. In Figure 7a), intuitive trends are observed within the 2-dimensional latent space. Specifically, samples with similar depths in the unseen test dataset are observed to be clustered together in the latent space, and a smooth transition between the shallow melt pools and deeper melt pools is also apparent. In Figure 7b), these trends are also shown to hold for the energy density of each melt pool, also demonstrating that the model is able to learn generalizable and intuitive representations of the underlying physical mechanisms of melt pool formation.
Figure 7: a) A t-SNE embedding plot based on the latent space embedding learned within the transformer directly before the prediction head, on the dataset of FLOW-3D simulations. A trend is observed where the samples with the largest depth are concentrated at the top right of the plot. b) A similar trend is observed with increasing energy density. c) The t-SNE embedding reveals that the model has the largest error at low energy density and depth values.
@@
Figure 8: a) A comparison of the Intersection-over-Union (IoU) metric as a function of the amount of data used for model training. These comparisons are made for both a model trained with randomly initialized weights, a model that has been transfer learned from FLOW-3D simulated melt pools, and a model that has been transfer learned from Eagar-Tsai simulated melt pools. b) A comparison of the correlation between the area extracted from the predicted cross-section to the area extracted from the ground truth cross-section, a function of the amount of data used for model training. These comparisons are made for both a model trained with randomly initialized weights, a a model that has been transfer learned from FLOW-3D simulated melt pools, and a model that has been transfer learned from Eagar-Tsai simulated melt pools.
We also create a dataset of low-fidelity simulations using the Eagar-Tsai model to simulate the melt pool. The Eagar-Tsai model has more restrictive assumptions than the high-fidelity simulation models. Specifically, the Eagar-Tsai model considers heat conduction as the only mode of heat dissipation, assumes that the thermal properties of the material remain constant with temperature, and neglects phase change phenomena [22]. Due to these assumptions, the melt pool temperatures computed are up to 10$\times$ higher than those reported in experiment [49]. However, the Eagar-Tsai model has a semi-analytical solution which is less expensive to compute than a FLOW-3D simulation at the same mesh resolution [56, 30]. To assess the performance of pre-training on this melt pool model for transfer learning, we create a dataset of 270 SS316L simulations and compute the surface temperature and depth contour with the same procedure described earlier for the FLOW-3D dataset. The semi-analytical equation for the Eagar-Tsai model is reproduced in Equation 4, where $A$ is the laser absorptivity, $\rho$ is the density of the material, $c_{p}$ is the heat capacity, $k$ is the thermal conductivity, $D$ is the thermal diffusivity defined as $D=\frac{k}{\rho c_{p}}$, $\sigma$ is the radius of the laser, $V$ is the velocity of the laser, and $P$ is the power of the laser. The parameters used to compute the solution at different power-velocity combinations are reported in C.
[…]
To apply the information learned by the model in simulation to the experimentally collected dataset, we first trained the model for 100 epochs on the simulation dataset as described earlier. Next, we fine-tune the trained model on a subset of the training partition of the single-track melt pool cross-sections for 10 epochs. We initialize the model with the weights learned from the simulation data training process, and continue the training process with a randomly selected sample of the training partition of the experimental dataset. Next, we evaluate the performance of the fine-tuned model on the original test partition of the dataset as the fraction of training data available as input varied. In Figure 8a), the correlation between the ground truth and predicted melt pool cross-sectional areas is shown to increase dramatically when a model pre-trained on simulation data is used for prediction, compared to a network with randomly initialized weights. We observe a greater increase in accuracy in the low dataset regime for the FLOW-3D pre-trained model when compared to the Eagar-Tsai pretrained model. However, as more data becomes available for the fine-tuning process, similar increases in accuracy are seen for both the FLOW-3D and Eagar-Tsai pre-trained models. A similar effect is observed in Figure 8b) for the IoU score of the predicted cross-sections. Therefore, fine-tuning a FLOW-3D pre-trained model on only 10 unique process parameter combinations enables predictions with an area correlation $R^{2}$ of 0.56, and increasing the amount of data to 20 unique process-parameter cross-sections enables prediction with an area correlation $R^{2}$ of 0.85.